OpenStudy (anonymous):

Explain how the graph of each given function is a transformation of the graph of y=x^2. The transformation is y=-2x^2. I know it is reflected over the x-axis, but how much does it stretch or shrink?

OpenStudy (anonymous):

Well, think about \( y = -2x^2 \) as the "reflection" of \( y = 2x^2 \) Now. \(y=2x^2\) is basically \( y= x^2 \) with all values doubled. This means, for every \(x\) you plug into \(x^2\) in \(2x^2\) you'll get a double value. This means \(2x^2\) is vertically stretched, as all values are doubled.. and \( y = -2x^2 \) is both vertically stretched and "reflected".

OpenStudy (anonymous):

I know it is stretched, but by how much compared to y=x^2?

OpenStudy (anonymous):

doubled

OpenStudy (anonymous):

Oh, so the answer would just be that the function is reflected over the x-axis and is stretched by double? Thanks!

OpenStudy (anonymous):

or stretched by factor of 2... Maybe you could compare with an example exercise to see how they want the answer to be like.

OpenStudy (anonymous):

Alright I will, thanks again!

OpenStudy (anonymous):

Sure np